# Simulating Longitudinal Data from Marginal Structural Models

### v1

@@ Abstract
Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and designing clinical trials. While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, marginal structural models require capturing causal effects that are marginal over time-dependent confounders, making it difficult to align conditional distributions with target marginal quantities. To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model. Recognizing the importance of time-to-event outcomes in clinical trials, we extend the method to accommodate survival models. Compared to existing approaches, our method offers several key advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids imposing restrictive assumptions on the data-generating process; and it is efficient as we need only evaluate analytic functions. This last benefit contrasts with methods that use computationally intensive techniques such as Monte Carlo approximations or numerical integration. Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility. Our method will allow researchers to effectively simulate data with target causal structures for their specific scenarios.
@@ 1 Introduction
Longitudinal data are very common in studies analyzing the causal effect of a treatment (or exposure) on an outcome, especially in the context of chronic diseases such as diabetes or HIV. However, treatments are rarely fixed for the entire period of observation, and their status often varies over time. Even in clinical trials with strict adherence protocols, intercurrent events—such as discontinuation of the prescribed treatment and use of additional medication—are inevitable. For example, participants in the treatment group might discontinue treatment due to side effects, while those in the control group could begin treatment or alternative therapies if their health declines. In observational studies, treatment initiation or cessation can occur at any time, contingent on changing circumstances. The ICH E9111International Council on Harmonization of Technical Requirements for Registration of Pharmaceuticals for Human Use [ICH] Statistical Principles for Clinical Trials [E9] emphasizes that intercurrent events must be explicitly considered when defining the clinical question of interest (European Medicines Agency, 2020). While the intention-to-treat (ITT) principle simplifies this to some extent by reducing time-varying treatment to a point-treatment problem, it may not always yield the most meaningful estimand from a clinical or regulatory perspective. Longitudinal data, however, allow us to more accurately capture the treatment dynamics, enabling estimation of causal quantities like the “on-treatment” estimand, which represents the treatment effect assuming treatment is administered as intended.
@@
A particular challenge in analyzing longitudinal data is time-varying confounding. This occurs when a time-dependent covariate predicts both the outcome and future treatment exposure, while past treatment influences the covariate. For example, in the case study by Robins et al. (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression. The presence of such time-varying confounding renders conventional adjustment methods, like regression and stratification, generally biased.
@@
Marginal structural models, first introduced by Robins (1997), have emerged as a powerful tool for estimating the causal effect of time-dependent exposures in the presence of time-varying confounding. Model parameters can be consistently estimated using inverse-probability-of-treatment-weighted estimators. Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), a systematic review of the literature by Clare et al. (2019) concluded that marginal structural models fitted with inverse probability of treatment weights are the most commonly used approach for longitudinal data with time-varying confounding due to their simplicity in both implementation and interpretation.
As novel causal inference methods for longitudinal data are continually proposed to address complex real-world scenarios, it is essential to evaluate their effectiveness for practical use. Simulation studies play a key role in this, by assessing important properties such as bias, confidence interval coverage, relative efficiency, and robustness to assumption violations. These evaluations guide statisticians in developing new methods and help analysts choose appropriate methods for specific research questions.
@@
Sample size calculation also benefits from simulations. Prior to conducting a trial or an observational study, it is crucial to determine the number of participants, or number of observational units required to achieve adequate statistical power for the hypothesis test (for example, superiority, inferiority or equivalence). An improperly sized sample is clearly problematic: if too few units are recruited it can lead to undetected non-zero effects; if too many are recruited this may be a waste of resources and potentially unethical if it means that more individuals are given an inferior treatment. While standard formulas can be used for time-fixed treatments, more complex scenarios with longitudinal data and time-varying treatments make deriving analytical formulas impractical or impossible. In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
@@
For both method evaluation and sample size calculation, simulating data that adheres to a known causal structure is crucial. However, simulating longitudinal data from a pre-specified marginal structural model is not a straightforward task. While generating data sequentially from structural equations—creating covariates, treatment status, and then outcome conditioned on past information—may seem intuitive, marginal structural models rely on marginal, rather than conditional models. This makes it difficult to sequentially simulate from conditional models in a way that aligns with a pre-specified marginal model; in particular, it is very difficult to sample from a null hypothesis.
@@
The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models. Section 3 covers the preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas. In Section 4, we introduce our proposed methodology and illustrate it in Section 5 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004). Section 6 extends the simulation approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, including an illustrative example. The paper concludes with a discussion in Section 7.
@@ 2 Related Literature
As noted in the previous section, the challenge lies in ensuring that simulated data adhere to a specified marginal model, rather than simply following sequential conditional distributions in a structural equation model. There are two possible strategies to address this issue:
@@
    Simulate data from the conditionals and then derive the corresponding marginal models. The g-formula, simplifying assumptions, and Monte Carlo approximations, can help bridge the gap between the simulated conditionals and the target marginal model.
@@
    Alternatively, bypass conditional simulations and develop methods to simulate directly from the target marginal model.
Both strategies are explored in existing approaches.
2.1 Simulate from the conditionals and derive the equivalent marginal models
Xiao et al. (2010) conducted a simulation study to evaluate the accuracy of two methods of fitting a Cox marginal structural model. They generated their dataset through the natural approach—sequentially through conditional models—and argued that under the assumption of a rare disease, i.e. a very low failure rate, the impact of non-collapsibility is negligible. In this scenario, the parameters of the underlying Cox marginal structural model may be approximated from the conditional model parameters. Notably, for this approximation to hold, there must be no confounders for the relationship between the time-varying confounder and the outcome, whether observed or unobserved. Essentially, their study did not directly simulate data from an marginal structural model but demonstrated a case where conditional simulation can serve as a reasonable approximation of such a model.
Young and Tchetgen Tchetgen (2014) explored bridging from conditionals to marginals using the g-formula in simulating from a discrete time Cox marginal structural model. In addition to standard identifying assumptions—consistency, positivity, and sequential exchangeability—they further assume that the g-formula uses standard parametric models, and hence derived the functional form of the marginal model. If the time-varying covariates consist only of discrete variables, then deriving the equivalent marginal relative hazard is theoretically possible, but it can be cumbersome. However, the authors noted that if the time-varying covariate includes continuous components, the summation in the above equation is replaced by integrals which may be challenging to evaluate due to the lack of closed-form solutions. The authors suggest potential simplifications such as the Markov property or approximation based on the rare disease assumption, as used by Xiao et al. (2010). A significant limitation arises when attempting to simulate data from a specific marginal structural model. Other than in very specific cases, such as for discrete or linear models, it is essentially impossible to reverse-engineer the series of equivalent conditional models. For instance, if we want to simulate from $h(t,\bar{a}_{t})=0.1+0.5a_{t}+0.3a_{t-1}-a_{t}a_{t-1}$, the corresponding series of conditionals is difficult to derive. This poses a drawback in real-world applications, especially for sample size determination, where experimenting with different effect sizes is necessary to generate power curves.
Sofrygin et al. (2017) developed the simcausal R package as a flexible, practical tool for simulating complex longitudinal data structures from non-parametric structural equation models. The package supports simulations that replicate realistic data scenarios common in causal inference, such as time-varying confounding and selection bias. Users can specify and simulate counterfactual data under various intervention regimes. A typical simcausal workflow involves parameterizing structural equations based on a causal directed acyclic graph and simulating data sequentially from the specified conditional distributions. Since deriving analytic expressions for marginal causal quantities, such as the average treatment effect, is often intractable with this type of conditional parameterization, users instead estimate these parameters through Monte Carlo approximations by simulating counterfactual data. While this approach is suitable for marginal causal quantities such as the average treatment effect, it may lack consistency for causal structures such as marginal structural models or conditional average treatment effect functions due to potential model misspecification. The authors acknowledged this limitation, as they allow users to specify a “working” model and cautioned that this approximation serves as the best parametric fit within the model class, without guaranteeing consistency.
Keogh et al. (2021) described how to simulate from an additive-hazard marginal structural model (Aalen, 1989). They utilized the result that if the conditional hazard model follows an additive form, the equivalent marginal hazard model will also be additive. Unlike Cox hazard ratios, the parameters of the additive hazard model are collapsible. Their strategy involved simulating data conditionally and then deriving the corresponding marginal hazard. However, while this approach ensures the correct form of the marginal model, the true values of the coefficients remain generally intractable. To address this, as with the workflow outlined in Sofrygin et al. (2017), the authors find a Monte Carlo approximation of the true parameters through simulating counterfactual data. A notable limitation of the additive hazard model is that it does not constrain the hazard to be non-negative. This becomes problematic when the hazard depends on continuous covariates, as ensuring positivity may be difficult in simulated data. Additionally, as with the approach by Young and Tchetgen Tchetgen (2014), this method does not allow direct specification of the model parameters, limiting its applicability in real-world scenarios.
2.2 Simulate directly from the target marginal model
Havercroft and Didelez (2012) approached this problem from a novel angle by simulating directly from the causal margin (in their case, a discrete Cox marginal structural model) and then introducing time-varying confounding to establish its relationships with the outcome and the treatment. Specifically, they used an inverse cumulative distribution function to simulate from the causal margin, with time-varying confounding introduced indirectly via a shared latent variable. This approach omits any direct edge from the time-varying confounder directly to the outcome. The primary limitation of this method is that time-varying confounding is not explicitly modeled but introduced via an artificial latent variable. As a result, the assumption that time-varying covariates have no direct effect on outcomes may not always hold in clinical contexts. For instance, in the HIV study example, this would assume that CD4 cell count has no direct effect on survival, which is unlikely. Nonetheless, their approach presents a valuable contribution.
Evans and Didelez (2024) recently proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly. This innovation enables direct simulation from the marginal structural model, addressing limitations identified in previous work. Specifically, it allows for explicit simulation of time-varying covariates and an exact marginal model. More details on the frugal parameterization are given in Section 3.1.
Building on this foundation, Seaman and Keogh (2024) applied the frugal parameterization to simulate longitudinal data with survival outcomes. Similar to the approach of Havercroft and Didelez (2012), they simulated the outcome using the inverse cumulative distribution function of the target causal margin. To handle high-dimensional time-varying covariates, they summarized their effects into a one-dimensional “risk score”, simplifying the dependence between covariates and the outcome to a bivariate copula. However, this approach has notable limitations. Since the cumulative distribution functions of the risk score are generally unknown, they relied on Monte Carlo approximations, which involve sampling a large number of individuals with the same treatment history and baseline variables. It is a computationally expensive task and must be performed at each time step. For example, with binary treatment, there could be up to $10242^{10}=$ 1024 combinations in terms of treatment history distribution functions to approximate at step 10 – potentially even more if baseline variables are included. The method becomes infeasible when treatment or baseline variables are continuous. The method also becomes impractical when treatment history or baseline variables have many possible combinations or are continuous. Moreover, the use of a risk score limits flexibility, as it prevents explicit modeling of specific covariate-outcome dependencies, such as the effect of age on survival. A more flexible approach would directly model these associations, making fuller use of the frugal parameterization.
The literature review highlights gaps in current methodologies and identifies key features desired in an improved simulation approach. These include the ability to bypass the conditional distribution of the outcome and simulate directly from the marginal structural model, robustness that avoids assumptions about the data-generating process, flexibility to accommodate diverse scenarios, and practicality for real-world applications.
@@ 3.1 The frugal parameterization
As discussed in the previous section, the fundamental challenge in simulating data from longitudinal marginal structural models arises from the fact that, while the joint data distribution can be easily factorized into conditionals for simulation, the causal quantity of interest is a marginal distribution. Earlier methods, as detailed in Section 2, tackle this by deriving the marginal distribution implied by the conditional factorization. However, Evans and Didelez (2024) significantly advanced this area with the introduction of the frugal parameterization, which explicitly centers the causal marginal and constructs the rest of the joint distribution around it. This breakthrough has made the simulation method presented in this paper possible.
@@
The first component, termed “the past”, refers to the distribution of covariates and treatment, which precede the outcome. The second component, the causal margin, aligns with the marginal structural model we aim to simulate, directly addressing the causal question at hand. The third component captures the dependency between the outcome $Y$ and the set of covariates $\bm{L}$ jointly. A common choice of this dependency, $\phi$, is a copula, which we will elaborate on in the following subsection.
@@ 3.2 Copulas
The term copula is derived from the Latin verb copulare, meaning “to join together.” We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure. This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$. According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\dots F_{d}$ can be expressed as
$F(x_{1},\dots,x_{d})=C\left\{F_{1}(x_{1}),\dots,F_{d}(x_{d})\right\}=C(u_{1},\dots,u_{d}),$
where $\bm{X}=(X_{1},\dots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\dots,d$. The copula then takes the form
$C(u_{1},\dots,u_{n})=F\{F_{1}^{-1}(u_{1}),\dots,F_{d}^{-1}(u_{d})\},\qquad(u_{1},\dots,u_{d})\in[0,1]^{d},$
where $x_{i}=F_{i}^{-1}(u_{i})$. The mathematical relationship between copula density function $c(u_{1},\dots,u_{d})$ and joint density function $f(x_{1},\dots,x_{d})$, applying the chain rule, is given by
$f(x_{1},\dots,x_{d})=c(u_{1},\dots u_{d})\prod_{i=1}^{d}f_{i}(x_{i}).$ (1)
@@ 3.2.1 Pair-copula construction
Building higher-dimensional copulas is generally recognized as a difficult problem. While there exist a wider variety of parametric bivariate copulas, the options for higher-dimensional copulas are limited, with Gaussian and Student t-copulas being common choices. Extending symmetric bivariate Archimedean copulas to multivariate settings is also possible but imposes strong restrictions, such as requiring the same level of dependence among all possible pairs. First introduced by Joe (1996), pair-copula constructions provide a flexible and intuitive approach to building multivariate copulas by modeling multivariate data as a series of bivariate copulas. This hierarchical structure incorporates additional variables at each level, using pair-copulas as the fundamental building blocks.
@@
where $\bm{x}$ is a $d$-dimensional vector; $x_{j}$ is an arbitrarily chosen component of $\bm{x}$, and $x_{-j}:=(x_{1},\ldots,x_{j-1},x_{j+1},\ldots,x_{d})$ denotes $\bm{x}$ with its $j$th entry removed. Applying (2) recursively to the conditional densities and, for convenience, choosing an order of $1,2,\dots d$ we have:
$f(y\,|\,\bm{x})$ $=c_{YX_{d}|\bm{X}_{1:{d-1}}}\{F(y\,|\,\bm{x}_{1:{d-1}}),F(x_{d}\,|\,\bm{x}_{1:{d-1}})\}$
$\quad\cdot c_{YX_{d-1}|\bm{X}_{1:{d-2}}}\{F(y\,|\,\bm{x}_{1:{d-2}}),F(x_{d-1}\,|\,\bm{x}_{1:{d-2}})\}$
[…]
$\quad\cdot c_{YX_{2}|X_{1}}\{F(y\,|\,x_{1}),F(x_{2}\,|\,x_{1})\}$
$\quad\cdot c_{YX_{1}}\{F(y),F(x_{1})\}\cdot f(y)$ (3)
$=f(y)\cdot\prod_{i=1}^{d}c_{YX_{d+1-i}|\bm{X}_{1:{d-i}}}\{F(y\,|\,\bm{x}_{1:{d-i}}),F(x_{d+1-i}\,|\,\bm{x}_{1:{d-i}})\},$ (4)
where each copula in the product in (4) is bivariate, i.e. a pair-copula. In principle, any ordering of the conditioned variables $x_{1},x_{2},\dots,x_{d}$ can be selected, although some orderings may be easier to use than others. In Section 4, we choose a particular ordering that is convenient and efficient for simulation.
In conclusion, this shows that under appropriate regularity conditions, a multivariate density can be expressed as a product of pair-copulas, acting on different conditional distributions.
3.2.2 Simulating from pair-copula-constructions

### v2

# Exact Simulation of Longitudinal Data from Marginal Structural Models

@@ Abstract
Simulating longitudinal data from specified marginal structural models is a crucial but challenging task for evaluating causal inference methods and clinical trial design. While data generation typically proceeds in a fully conditional manner using structural equations according to a temporal ordering, it is difficult to ensure alignment between conditional distributions and the target marginal causal effects. This misalignment presents a fundamental challenge in simulating data that adheres to marginal structural model specifications. To address this, we propose a flexible and efficient algorithm for simulating longitudinal data that adheres exactly to a specified marginal structural model. Recognizing the importance of time-to-event outcomes in clinical research, we extend our approach to accommodate survival models. Compared to existing approaches, our method offers several advantages: it enables exact simulation from a known causal model rather than relying on approximations; avoids restrictive assumptions about the data-generating process; and remains computationally efficient by requiring only the evaluation of analytic functions. This last benefit contrasts with methods that use computationally intensive techniques such as Monte Carlo approximations or numerical integration. Through simulation studies replicating realistic scenarios, we validate the method’s accuracy and utility. Our method will facilitate researchers in effectively simulating data with target causal structures for their specific scenarios.
@@ 1 Introduction
Longitudinal data are fundamental for studies analyzing the causal effect of a treatment or exposure on an outcome, especially for research into chronic diseases such as diabetes or HIV. Unlike cross-sectional studies, longitudinal designs capture the dynamic nature of treatment regimens, which often vary over time due to clinical decisions, patient adherence, or external factors. Even in clinical trials with strict adherence protocols, intercurrent events—such as treatment discontinuation or supplementary medication use—are common. By accounting for these treatment dynamics, longitudinal data enable the estimation of meaningful causal quantities, such as the “on-treatment” estimand, which quantifies the treatment effect assuming treatment is administered as intended.
@@
A central challenge in analyzing longitudinal data is addressing time-varying confounding, where a time-dependent covariate predicts both the outcome and future treatments, while itself being influenced by past treatments. For example, in the case study in Robins et al. (2000) on the effect of zidovudine (AZT) treatment on HIV infection, the biomarker CD4 lymphocyte count is both affected by prior AZT treatment and is predictive of future therapy initiation and HIV progression. The presence of such time-varying confounding renders conventional adjustment methods, like regression and stratification, generally biased.
@@
Marginal structural models, introduced by Robins (1997), have emerged as a powerful tool for estimating causal effects under time-varying confounding. Model parameters can be consistently estimated using inverse probability of treatment weighting (Robins, 1999, 2000). Among alternative methods, including structural nested models estimated through g-computation (Robins, 1989; Robins and Tsiatis, 1991), marginal structural models have gained prominence due to their interpretability and ease of implementation. A systematic review by Clare et al. (2019) highlights their widespread use in applied research.
@@
As causal inference methods for longitudinal data are continually proposed to address complex real-world settings, simulation studies are essential to evaluate their effectiveness for practical use. These studies assess important properties such as bias, confidence interval coverage, relative efficiency, and robustness to violations of the assumptions used, guiding both methodological development and application. Sample size calculations also benefit from simulations. Prior to conducting a trial or an observational study, it is crucial to determine the number of participants or number of observational units required to achieve adequate statistical power. While standard formulas can be used for time-fixed treatments, more complex scenarios with time-varying treatments make deriving analytical formulas impractical or impossible. In such cases, simulation is invaluable, allowing researchers to estimate sample sizes based on realistic data-generating processes and causal structures.
@@
A fundamental requirement for both method evaluation and sample size determination is the ability to simulate data that conform to a specified causal model. However, simulating longitudinal data from a pre-specified marginal structural model is not a straightforward task. While generating data sequentially from structural equations—creating covariates, treatment status, and then the outcome conditioned on past information—may seem intuitive, these models rely on conditional rather than marginal specifications. This makes it difficult to sequentially simulate with conditional distributions in a way that aligns with a pre-specified marginal model; in particular, it is extremely challenging to sample from a null hypothesis —the so-called g-null paradox (Robins and Wasserman, 1997).
@@
The remainder of this paper is structured as follows: Section 2 reviews existing methods for simulating longitudinal data from causal models. Section 3 introduces key preliminaries, including the frugal parameterization (Evans and Didelez, 2024) and copulas. In Section 4, we introduce our proposed methodology and illustrate it in Section 4.4 through a simulation study replicating a real-world study in epidemiology (Bodnar et al., 2004). Section 5 extends the simulation approach to time-to-event data, enabling simulations from survival models such as Cox marginal structural models, including an illustrative example. The paper concludes with a discussion in Section 6. Example code for the simulations in Sections 4.4 and 5.3 is available at https://github.com/XiLinStats/LongSiMSM.
@@ 2 Related Literature
As noted in Section 1, the challenge lies in ensuring that the simulated data adhere to a specified marginal model, rather than simply following sequential conditional distributions in a structural equation model. Existing approaches address this issue through two strategies:
@@
    Simulating from conditional models and deriving or approximating the corresponding marginal models post-hoc.
@@
    Bypassing conditional simulations and simulating directly from the target marginal model.
Several studies have explored the first approach. Xiao et al. (2010) used sequential conditional models to approximate Cox marginal structural models under the rare disease assumption, i.e. very low failure rates, which mitigates the impact of noncollapsibility but assumes the absence of unmeasured confounding between time-varying confounders and outcomes. Young and Tchetgen Tchetgen (2014) employed the g-formula to bridge conditional and marginal models in discrete-time Cox marginal structural models. While theoretically feasible for discrete time-varying covariates, deriving marginal models becomes intractable for continuous covariates due to the need for numerical integration. Simplifications, such as the Markov property or rare disease assumptions, are often necessary but limit generalizability. Sofrygin et al. (2017) developed the simcausal R package for flexible simulation from non-parametric structural equation models. The workflow involves specifying structural equations and sequentially simulating data from conditional distributions. However, deriving analytic expressions for marginal causal quantities under this conditional parameterization is often intractable and prone to model misspecification. To address this, users typically specify low-dimensional working marginal structural models (Neugebauer and van der Laan, 2007) and estimate their parameters via Monte Carlo simulations of counterfactual data. This introduces two limitations: first, the true marginal model remains unknown; second, reverse-engineering simulations to match a target causal function is challenging. Keogh et al. (2021) described simulation from additive-hazard marginal structural models (Aalen, 1989). They utilized the result that if the conditional hazard model follows an additive form, the equivalent marginal hazard model will also be additive. While the marginal model retains the correct form, true parameter values are generally intractable and require Monte Carlo approximations. A further limitation is the difficulty in ensuring non-negative hazards for continuous covariates.
Alternative methods focus on simulating directly from the target marginal model. Havercroft and Didelez (2012) simulated directly from a discrete Cox marginal structural model, introducing time-varying confounding indirectly via a shared latent variable. However, this approach assumes no direct effect of time-varying confounders on outcomes, which may not hold in practice, e.g. the effect of CD4 count on HIV survival (Robins et al., 2000). Evans and Didelez (2024) proposed a frugal parameterization that prioritizes the causal margin and parameterizes the surrounding variables accordingly, enabling direct simulation from the exact marginal model while explicitly modeling time-varying covariates. This addresses key limitations of earlier methods. Building on this, Seaman and Keogh (2024) applied the frugal parameterization to simulate survival outcomes, using a one-dimensional “risk score” to summarize multi-dimensional covariates. While simplifying dependence structures, this method relies on computationally expensive Monte Carlo approximations to estimate the cumulative distribution functions of the risk score, making it infeasible for continuous treatments or baseline variables.
The aims for a new simulation approach therefore include the ability to: precisely specify the marginal structural model; include multiple time-varying covariates, treatments, outcomes, and censoring; have good control over their distributions; avoid restrictive assumptions about the data-generating process. In addition, we would like it to scale well in the number of time-points and variables, and be applicable to real-world problems.
@@ 3.1 The frugal parameterization
As discussed in the previous section, the fundamental challenge in simulating data from longitudinal marginal structural models arises from the fact that, while the joint data distribution can be easily factorized into conditionals for simulation, the causal quantity of interest is a marginal distribution. Earlier methods, as detailed in Section 2, tackled this by deriving the marginal distribution implied by the conditional factorization. However, Evans and Didelez (2024) advanced this area with the introduction of the frugal parameterization, which explicitly parameterizes the causal marginal and constructs the rest of the joint distribution around it. This development has made the simulation method presented in this paper possible.
@@
A common choice of this dependency measure is a copula, which we elaborate on in the following subsection.
@@ 3.2 Copulas
We can think of continuous multivariate distributions as the joining together of the univariate marginals (or margins) and the dependence structure. This structure can be represented by a copula, which is a multivariate distribution function $C:[0,1]^{d}\rightarrow[0,1]$ with margins uniformly distributed on $(0,1)$. According to Sklar’s theorem (Sklar, 1959, 1973), any continuous multivariate distribution $F$ with marginals $F_{1},\ldots,F_{d}$ can be expressed as
$F(x_{1},\ldots,x_{d})=C\left\{F_{1}(x_{1}),\ldots,F_{d}(x_{d})\right\}=C(u_{1},\ldots,u_{d}),$
where $\bm{X}=(X_{1},\ldots,X_{d})$ is a vector of random variables, and $u_{i}=F_{i}(x_{i})$, $i=1,\ldots,d$. The copula then takes the form
$C(u_{1},\ldots,u_{d})=F\{F_{1}^{-1}(u_{1}),\ldots,F_{d}^{-1}(u_{d})\},\qquad(u_{1},\ldots,u_{d})\in[0,1]^{d}.$
@@ 3.2.1 Pair-copula construction
While there exist a wide variety of parametric bivariate copulas, the options for higher-dimensional copulas are limited, with Gaussian and Student t-copulas being common choices. Extending symmetric bivariate Archimedean copulas to multivariate settings is also possible but imposes strong restrictions, such as requiring the same level of dependence among all possible pairs. Building on the work of Joe (1996), Bedford and Cooke (2001, 2002) introduced pair-copula constructions, a flexible and intuitive approach to constructing multivariate copulas by modeling multivariate data as a series of bivariate copulas. This hierarchical structure incorporates additional variables at each level, using pair-copulas as the fundamental building blocks.
@@
where $\bm{x}$ is a $d$-dimensional vector; $x_{j}$ is an arbitrarily chosen component of $\bm{x}$, and $x_{-j}:=(x_{1},\ldots,x_{j-1},x_{j+1},\ldots,x_{d})$ denotes $\bm{x}$ with its $j$th entry removed. Applying (1) recursively to the conditional densities, we have:
$f(y\,|\,\bm{x})$ $=f(y)\cdot\prod_{i=1}^{d}c_{YX_{d+1-i}|\bm{X}_{1:{d-i}}}\{F(y\,|\,\bm{x}_{1:{d-i}}),F(x_{d+1-i}\,|\,\bm{x}_{1:{d-i}})\},$ (2)
where each copula in the product in (2) is bivariate; this defines a pair-copula. In principle, any ordering of the conditioned variables $x_{1},x_{2},\ldots,x_{d}$ can be selected, although some orderings may be easier to use than others. In our proposed method, we choose an ordering that is convenient and efficient for simulation.
In conclusion, under appropriate regularity conditions, a multivariate density can be expressed as a product of pair-copulas, acting on different conditional distributions.
3.2.2 Simulating from pair-copula constructions